# Fear over Friends: Examining the Perceived Influence of Others on Vaccination Decisions

### v1

@@ Abstract
In this study, we examine the perceived influence of others, across both strong and weak social ties, on vaccination decisions in the United States. In particular, we focus on the case study of COVID-19 vaccinations. We add context to social influence by measuring related concepts, such as perceived agreement of others and perceived danger of COVID-19 to others. We find that vaccinated populations perceived more influence from their social circles than unvaccinated populations. This finding holds true across various social groups, including family, close friends, co-workers, and neighbors. Indirect measures of social influence also followed this trend. Vaccinated participants perceived COVID-19 as more dangerous to their social circles than unvaccinated participants and perceived that others agreed with their decision to get vaccinated more than unvaccinated participants perceived others to agree with their decision to not get vaccinated. Despite the clear differences in perceived social influence across the groups, we find through open-ended responses that both vaccinated and unvaccinated participants frequently cited fear as a motivating factor in their decision, rather than social influence: vaccinated participants feared COVID-19, while unvaccinated participants feared the vaccine itself.
@@ 1 Introduction
One thing learned from COVID-19 is that the spread of a pandemic in nations around the world depended on not only on accurate epidemiological information and government responses Prather_etal_2020 ; Aksoy_etal_2020 ; Hale_etal_2020 ; Bedford_etal_2020 ; Frey_etal_2020 ; Munster_etal_2020 ; Chowell&Mizumoto_2020 ; Zhang&Qian_2020 but also the varied behaviors of individuals, groups, and cultures Maharaj&Kleczkowski_2012 ; Funk_etal_2010 ; Guiteras_etal_2015 ; Zhang&Centola_2019 ; gharzai2020playing ; Ruck_etal_2021a . Since culture is the context for behavior Zhang&Centola_2019 ; Gelfand_etal_2021 , and cultural values vary substantially across the world Aksoy_etal_2020 ; Inglehart&Welzel_2005 ; Ruck_etal_2020 , the effect of COVID-19 in terms of cases and deaths, was affected not only by government actions but by cultural values in their populations Ruck_etal_2021a .
@@
These varied behaviors have been well studied, particularly early in the COVID-19 pandemic, before vaccines were widely available. These studies have used socioeconomic and public health variables to explain COVID-19 variation within the United States Desmet&Wacziarg_2020 ; Ruck_etal_2021b and also globally deOliveira_etal_2020 . In the context of the literature on the effect of socioeconomic factors on COVID-19 Aksoy_etal_2020 ; Hale_etal_2020 ; Bedford_etal_2020 ; Frey_etal_2020 ; Munster_etal_2020 ; Chowell&Mizumoto_2020 ; Zhang&Qian_2020 ; Maharaj&Kleczkowski_2012 ; Funk_etal_2010 ; Guiteras_etal_2015 ; Zhang&Centola_2019 , cultural effects have been examined in concert with known risks such as obesity and advanced age, together with variables describing government efficiency and public trust in institutions Ruck_etal_2021a ; Ruck_etal_2021b .
@@
In terms of the behavioral dynamics of vaccination, those who are not vaccinated may feel less urgency due to a lowered perceived risk of infection verelst2016 ; fineberg2013paradox ; jansen2003measles . Furthermore, even those infected with a coronavirus or flu, may not exhibit obvious symptoms to others Li_etal_2020 and inadequate testing of populations Munster_etal_2020 can mean that the majority of infections in a population are undocumented Li_etal_2020 ; Munster_etal_2020 ; Bendavid_etal_2020 ; Baud_etal_2020 ; Marchant_etal_2020 . This can lead individuals to underestimate the risk of a virus, especially in lieu of more overt socio-economic concerns in people’s daily lives Mani_etal_2013 ; Trueblood_etal_2020 . Social influence from peers may be weak because—unlike visible protections such as mask-wearing or conspicuous lack of people in public spaces—vaccination is often comparatively less visible. In this situation, the perceived benefits of vaccination may be outweighed by economic and/or psychological costs Trueblood_etal_2020 ; Maharaj&Kleczkowski_2012 ; Pritchard_etal_2022a ; greenstone2020does .
@@
The complexity of the levels of voluntary vaccination lies in the multiple drivers of behavioral change, including information and social learning Carrignon_etal_2022 ; Funk_etal_2010 ; Guiteras_etal_2015 ; Christakis&Fowler_2013 ; Bentley_etal_2011 ; Silk_etal_2022 . Ideally, decisions would be determined by their intrinsic payoffs within their socio-ecological environment Gibson&Lawson_2015 . In the real world, decisions are made by people who combine observational learning, which produces noisy information, and social learning, which diffuses that information to others Hoppitt&Laland_2013 ; Rendell_etal_2011 . The transparency of learning is a crucial parameter Hauser_etal_2014 ; the less transparent the payoffs are, the more noise and heterogeneity enters into the behavioral dynamics Li&Lee_2009 ; Caiado_etal_2016 ; Vidiella_etal_2022 . While humans are motivated to avoid cues associated with pathogens, through emotions such as fear or disgust Schrock_etal_2020 , COVID-19 and similar viruses display few cues in asymptomatic cases kronbichler2020asymptomatic .
@@
These considerations suggest that drivers of social distancing include two dynamic factors: observable risks and observable behaviors. In everyday life, we may expect more intimate social influence, such as advice from parents or other close family members, to be the strongest for vaccinations. Vaccination can be made more visible of course through public outreach, but its acceptance will vary based on group norms and potentially large differences in risk perception due to misinformation. Acceptance may also vary based on a range of cultural memory effects, such as the past misuse of vaccination on marginalized groups. In contrast, if genuine benefits of vaccination become more transparent to the public, individual cost-benefit decisions can support the behavior spreading through social influence Carrignon_etal_2022 ; Guiteras_etal_2015 ; Silk_etal_2021 ; Pritchard_etal_2022b ; Banerjee_etal_2013 .
@@
Here we examine one facet of this complex problem: the perceived social influence of others on vaccination decisions. While prior work has shown that social influence matters in an array of health-related decisions, including social distancing and mask wearing during the COVID-19 pandemic tunccgencc2021social , examining social influence on COVID-19 vaccine decisions has had comparably less study, particularly across strong and weak social ties. Hence, our work seeks to fill this gap and add robustness to the current findings in the literature. More broadly, further work in this area adds to our understanding of the interactions between socio-cultural dynamics and health-related decision making.
Our four research questions are as follows:
@@
  • RQ4: If not by social influence, what other reasons do vaccinated and unvaccinated populations think influence their vaccination decision?
@@ 2 Methods
To answer our core research questions, we designed and deployed a survey on the online survey platform Prolific (prolific.co). The survey contained questions on demographics, perceived influence of varying social circles, perceived agreement of varying social circles, perceived danger of COVID-19 to varying social circles, and an open-ended reasoning question.
@@
We provide details about each type of question below.
First, we collected several types of demographic information for comparison to previous studies on COVID-19 vaccination and related behaviors. These demographic questions included:
  • Political leaning - We asked “What is your political leaning?” . Participants could answer on a 5-point scale: very liberal, liberal, moderate, conservative, or very conservative.
  • Rural-urban residency - We asked participants to provide the U.S. zip code in which they currently reside. This question allowed us to map participants to Rural-Urban Continuum Codes from the U.S. Department of Agriculture [^1] https://www.ers.usda.gov/data-products/rural-urban-continuum-codes.aspx. Rural-Urban Continuum Codes range from 1 - counties in metro areas of 1 million population or more - to 9 - counties that are completely rural or less than 2,500 urban population, not adjacent to a metro area.
  • Rural-urban identity - Following the questioning used in lunz2022rural , we asked “Regardless of where you currently live, do you usually think of yourself as a city person, a suburban person, a small-town person, a rural or country person, or something else?”, giving us a 4-point scale from country person to city person with something else as an outlier. According to lunz2022rural , rural social identification - “a psychological attachment to being from a rural area or small town” - is different than “simply living in a rural area”. Lunz Trujillo (2022) demonstrated this by showing that rural social identification predicts greater anti-intellectualism than rural residency alone. Hence, in our study we wanted to capture this psychological trait in addition to where one lives.
  • Education - We asked “What is your level of education?”. Participants could answer on a 7-point scale: less than high school, high school graduate, some college, 2 year degree, 4 year degree, professional degree, doctorate.
  • Income - We asked “What is your approximate household income?”. Participants could answer on a 12-point scale: less than \$10,000 to more than \$150,000.
  • Gender - We asked “How do you describe yourself?”. Participants could answer with one of four answers: male, female, non-binary/third gender, prefer not to say.
  • Race - We asked “What is your racial ethnicity?”. Participants could answer with one of seven answers: white, black, Asian, Hispanic or Latino, native American or Alaskan, Native Hawaiian or Pacific Islander, or other
Next, we asked participants “How much influence did each social group have on your decision to get vaccinated?” or “ How much influence did each social group have on your decision to not get vaccinated?” as a matrix Likert Scale question across five groups: family members, close friends, co-workers, church or social club, and neighbors or community members. Participants could answer on a 5-point scale: none at all, a little, a moderate amount, a lot, or a great deal.
To add context to our question about social influence, we also asked participants about social agreement with their vaccination decision and their perceived danger of COVID-19 to others. Specifically, we asked participants “How much do you think each social group agrees with your decision to get vaccinated?” or “How much do you think each social group agrees with your decision to not get vaccinated?” as a matrix Likert Scale question across five groups: family members, close friends, co-workers, church or social club, and neighbors or community members. Participants could answer on a 3-point scale: disagree, do not care, or agree. We also asked participants “How likely do you think members of each social group are to get severe COVID-19 or die from COVID-19?” as a matrix Likert Scale question across five groups: family members, close friends, co-workers, church or social club, and neighbors or community members. Participants could answer on a 5-point scale: extremely unlikely, somewhat unlikely, neither likely nor unlikely, somewhat likely, or extremely likely.
Next, we wanted to understand both the size of participants social circles and how much they trust members of their social circles. Hence, we asked participants “To the best of your ability, type in the first names or relationship titles (mother, best friend, cousin, co-worker, etc.) of each person you voluntarily had a conversation with in the past week.” and “Of those people you listed in the previous question, how many of them would you turn to for advice with a major personal problem?” to approximate the number of people they interact with on a regular basis and how much they trust those individuals. This line of questioning is similar to the questions asked in prior work to approximate the size of one’s social circle dunbar1995social ; tunccgencc2021social .
Lastly, we asked “In your own words, why did you get a COVID-19 vaccine?” or “In your own words, why did you not get a COVID-19 vaccine?” as an open-ended question to elicit reasoning outside of our social influence questions. Our goal with this question is to capture if participants primary reasoning for getting or not getting vaccinated is indeed based on social cues or if it is based on something else.
@@
In the data analyzed, we use data from both the pilot deployment and actual deployment, totalling to 1015 responses. After filtering out responses that did not pass our attention check question, we had 1000 survey responses to analyze between the three deployments, with 486 unvaccinated participants and 514 vaccinated participants.
To analyze our ordered Likert Scale questions, we leverage ordinal logistic regression. Each model uses an ordered categorical question as the dependent variable and if the participant was in the vaccinated group or unvaccinated group has the independent variable. This method allows us to estimate changes in units of ordered logits between the two groups and if those changes are statistically significant. We describe this process in more detail in Section 3 and Table 1.
@@
To analyze our open-ended reasoning question, we take a multi-round, open-coding approach. First, two of the authors individually went over the open-ended responses from the vaccinated group and the unvaccinated group to identify broad codes, continuing until the list of codes was stable. The two authors then discuss their codes to develop a combined list of codes, each with specific definitions of what reasoning fits into the category. The final set of codes can be found in Tables 3 and 4. Then the same two authors coded all of the responses for both groups (vaccinated and not). Since the respondents could describe as many reasons as they wanted, each participants response could fall into more than one category. After the second round of coding, agreement was computed, where a response is said to have agreement if all codes match across the two coders. In the unvaccinated group, 79.84% of the responses had full agreement, while the vaccinated group, 85.38% of the responses had full agreement. The two authors met again to discuss and resolve these conflicts. Disagreements in the unvaccinated group were mostly due to interpreting what entity was being distrusted (science, government, media, pharmaceutical companies, general), hence we merged these to a broader category of Distrust. For example, some response were very specific: “Because there’s not enough research to determine if it’s safe and I do not trust the government or media.” while others were more general: “i dont trust any of it”. Disagreements in the vaccinated group mostly were due to differentiating the categories Protect others versus Right thing to do. In this case, we opted for stricter definitions of each to avoid misinterpretation. For example, a response had to use the phrase “it was the right thing to do” to be included in the category Right thing to do, while calls to protecting others, such as “To protect my family, community” or “I wanted to protect others from catching Covid and help reduce the spread” were put in the category Protect others.
Definitions and examples of the final agreed upon results can be found in Tables 2, 3, and 4.
@@ 3 Results
Figure 1: U.S. counties where at least one survey participant resides. Participants in our survey lived in 882 unique U.S. counties, with 471 counties represented in the vaccinated group and 461 counties represented in the unvaccinated group.
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[…]
figureDistributions of (a) political leaning, (b) rural-urban identity, and (c) education of vaccinated and unvaccinated participants.

### v2

# Perceived Social Influence on Vaccination Decisions: A COVID-19 Case Study

@@ Abstract
In this study, we examine the perceived influence of others, across both strong and weak social ties, on COVID-19 vaccination decisions in the United States. We add context to social influence by measuring related concepts, such as perceived agreement of others and perceived danger of COVID-19 to others. We find that vaccinated populations perceived more influence from their social circles than unvaccinated populations. This finding holds true across various social groups, including family, close friends, and neighbors. Vaccinated participants perceived that others agreed with their decision to get vaccinated more than unvaccinated participants perceived others to agree with their decision to not get vaccinated. Despite the clear differences in perceived social influence and agreement across the groups, the majority of participants across both vaccinated and unvaccinated populations perceived no social influence from all social group in their decisions. Aligning with this result, we find through open-ended responses that both vaccinated and unvaccinated participants frequently cited fear as a motivating factor in their decision, rather than social influence: vaccinated participants feared COVID-19, while unvaccinated participants feared the vaccine itself.
@@ 1 Introduction
One thing learned from COVID-19 is that the spread of a pandemic in nations around the world depended on not only on accurate epidemiological information and government responses Prather_etal_2020 ; Aksoy_etal_2020 ; Hale_etal_2020 ; Bedford_etal_2020 ; Frey_etal_2020 ; Munster_etal_2020 ; Chowell&Mizumoto_2020 ; Zhang&Qian_2020 but also the varied behaviors of individuals, groups, and cultures Maharaj&Kleczkowski_2012 ; Funk_etal_2010 ; Guiteras_etal_2015 ; Zhang&Centola_2019 ; gharzai2020playing ; Ruck_etal_2021a . Since culture is the context for behavior Zhang&Centola_2019 ; Gelfand_etal_2021 , and cultural values vary substantially across the world Aksoy_etal_2020 ; Inglehart&Welzel_2005 ; Ruck_etal_2020 , the effect of COVID-19 in terms of cases and deaths, was affected not only by government actions but by cultural values and social norms in their populations Ruck_etal_2021a ; moehring2021surfacing .
@@
These varied behaviors have been well-studied, particularly early in the COVID-19 pandemic, before vaccines were widely available. These studies have used socioeconomic and public health variables to explain COVID-19 variation within the United States Desmet&Wacziarg_2020 ; Ruck_etal_2021b and also globally deOliveira_etal_2020 . In the context of the literature on the effect of socioeconomic factors on COVID-19 Aksoy_etal_2020 ; Hale_etal_2020 ; Bedford_etal_2020 ; Frey_etal_2020 ; Munster_etal_2020 ; Chowell&Mizumoto_2020 ; Zhang&Qian_2020 ; Maharaj&Kleczkowski_2012 ; Funk_etal_2010 ; Guiteras_etal_2015 ; Zhang&Centola_2019 , cultural effects have been examined in concert with known risks such as obesity and advanced age, together with variables describing government efficiency and public trust in institutions Ruck_etal_2021a ; Ruck_etal_2021b .
@@
In terms of the behavioral dynamics of vaccination, those who are not vaccinated may feel less urgency due to a lowered perceived risk of infection verelst2016 ; fineberg2013paradox ; jansen2003measles . Furthermore, even those infected with a coronavirus or flu may not exhibit obvious symptoms to others Li_etal_2020 . Inadequate testing of populations Munster_etal_2020 can mean that the majority of infections in a population are undocumented Li_etal_2020 ; Munster_etal_2020 ; Bendavid_etal_2020 ; Baud_etal_2020 ; Marchant_etal_2020 . This can lead individuals to underestimate the risk of a virus, especially in context with other socio-economic concerns Mani_etal_2013 ; Trueblood_etal_2020 . Social influence from peers to vaccinate may be weak because—unlike visible protections such as mask-wearing or conspicuous lack of people in public spaces—vaccination is often comparatively less visible. In this situation, the perceived benefits of vaccination may be outweighed by economic and/or psychological costs Trueblood_etal_2020 ; Maharaj&Kleczkowski_2012 ; Pritchard_etal_2022a ; greenstone2020does .
@@
The complexity of the levels of voluntary vaccination lies in the multiple drivers of behavioral change, including information and social learning Carrignon_etal_2022 ; Funk_etal_2010 ; Guiteras_etal_2015 ; Christakis&Fowler_2013 ; Bentley_etal_2011 ; Silk_etal_2022 . Ideally, decisions would be determined by their intrinsic payoffs within their socio-ecological environment Gibson&Lawson_2015 . In the real world, decisions are made by people who combine observational learning, which produces noisy information, and social learning, which diffuses that information to others Hoppitt&Laland_2013 ; Rendell_etal_2011 . The transparency of learning is a crucial parameter Hauser_etal_2014 ; the less transparent the payoffs are, the more noise and heterogeneity enters into the behavioral dynamics Li&Lee_2009 ; Caiado_etal_2016 ; Vidiella_etal_2022 . While humans are motivated to avoid cues associated with pathogens, through emotions such as fear or disgust Schrock_etal_2020 , COVID-19 and similar viruses displayed few cues in asymptomatic cases kronbichler2020asymptomatic .
@@
These considerations suggest that drivers of social distancing include two dynamic factors: observable risks and observable behaviors. In everyday life, we may expect more intimate social influence, such as advice from parents or other close family members, to be the strongest for vaccinations. Vaccination can be made more visible of course through public outreach, but its acceptance will vary based on group norms and potentially large differences in risk perception due to misinformation. Acceptance may also vary based on a range of cultural memory effects, such as the past misuse of vaccination on marginalized groups. In contrast, if genuine benefits of vaccination become more transparent to the public, individual cost-benefit decisions can support the behavior spreading through social influence Carrignon_etal_2022 ; Guiteras_etal_2015 ; Silk_etal_2021 ; Pritchard_etal_2022b ; Banerjee_etal_2013 ; bentley2024cultural .
@@
Here we examine one under-researched facet of this complex problem: the perceived social influence of others on vaccination decisions. While prior work has shown that social influence matters in an array of health-related decisions, including social distancing and mask wearing during the COVID-19 pandemic tunccgencc2021social , examining social influence on COVID-19 vaccine decisions has had comparably less study, particularly across strong and weak social ties. Hence, our work seeks to fill this gap and add robustness to the current findings in the literature. More broadly, further work in this area adds to our understanding of the interactions between socio-cultural dynamics and health-related decision making. Our three research questions are as follows:
@@
From our quantitative analysis, we find that vaccinated people perceived significantly more social influence from their families, close friends, and neighbors in their decision to get vaccinated than unvaccianted people perceived in their decision to not get vaccinated. However, notably, the majority of participants across both vaccinated and unvaccinated populations perceived no social influence from all social group. Likewise, vaccinated people perceived significantly more agreement with their decision to get vaccinated from their families, close friends, coworkers, and neighbors than unvaccinated people perceived with decision to not get vaccinated. Surprisingly, perceived danger of COVID-19 to varying social circles did not significantly differ across vaccinated and unvaccinated populations. Yet, through open-ended responses that both vaccinated and unvaccinated participants frequently cited fear as a motivating factor in their decision, rather than social influence or agreement: vaccinated people feared COVID-19, while unvaccinated people feared the vaccine itself.
@@ 2 Methods
2.1 Survey
To answer our research questions, we designed and deployed a survey on the online survey platform Prolific (prolific.co). The survey contained questions on perceived influence of varying social circles, perceived agreement of varying social circles, perceived danger of COVID-19 to varying social circles, and an open-ended reasoning question.
Specifically, we asked participants “How much influence did each social group have on your decision to get vaccinated?” or “ How much influence did each social group have on your decision to not get vaccinated?” as a matrix Likert Scale question across five groups: family members, close friends, co-workers, church or social club, and neighbors or community members. Participants could answer on a 5-point scale: none at all, a little, a moderate amount, a lot, or a great deal. Moreover, we asked participants “How much do you think each social group agrees with your decision to get vaccinated?” or “How much do you think each social group agrees with your decision to not get vaccinated?” as a matrix Likert Scale question across the same five groups. Participants could answer on a 3-point scale: disagree, do not care, or agree. We also asked participants “How likely do you think members of each social group are to get severe COVID-19 or die from COVID-19?” as a matrix Likert Scale question across the same five groups. Participants could answer on a 5-point scale: extremely unlikely, somewhat unlikely, neither likely nor unlikely, somewhat likely, or extremely likely.
To add context to these core questions, we wanted to understand both the size of participants social circles and how much they trust members of their social circles. Hence, we asked participants “To the best of your ability, type in the first names or relationship titles (mother, best friend, cousin, co-worker, etc.) of each person you voluntarily had a conversation with in the past week.” and “Of those people you listed in the previous question, how many of them would you turn to for advice with a major personal problem?” to approximate the number of people they interact with on a regular basis and how much they trust those individuals. This line of questioning is similar to the questions asked in prior work to approximate the size of one’s social circle dunbar1995social ; tunccgencc2021social .
Lastly, we asked “In your own words, why did you get a COVID-19 vaccine?” or “In your own words, why did you not get a COVID-19 vaccine?” as an open-ended question to elicit reasoning outside of our social influence questions. Our goal with this question was to add further context, as participants’ primary reasoning for getting or not getting vaccinated may not be based on social cues.
In addition, we collected several types of demographic information for comparison to previous studies on COVID-19 vaccination and related behaviors. These demographic questions included: gender, race, political leaning, income, education, rural-urban residency, and rural-urban identity (as defined by lunz2022rural ). Details on each question and the scales used can be found in the supplemental material.
Deployment
@@ 2.1 Survey
2.2 Analysis
@@
We analyzed the data from both the pilot deployment and actual deployment, totalling to 1015 responses. After filtering out responses that did not pass our attention check question, we had 1000 survey responses to analyze between the three deployments, with 486 unvaccinated participants and 514 vaccinated participants.
Quantitative
To analyze this data, we fit four ordinal logistic regression models, one for demographics alone and one for each of our three research question, where the dependent variable was a binary categorical variable representing the group a participant belonged to (0 for unvaccinated, 1 for vaccinated). Our independent variables in these models were our key demographic variables and the influence, agreement, or danger perceptions across the five social circles. These models allowed us to compare the predictive power of perceived influence, agreement or danger relative to demographics and across strong and weak ties. All the independent variables were coded as ordered scales starting at 1. For example, for Family Influence, ‘None at all’ is coded as 1, ‘A little’ is coded as 2, ‘A moderate amount’ is coded as 3, ‘A lot’ is coded as 4, ‘A great deal’ is coded as 5. Hence, the Likert scale value increases as Family Influence increases.. The coefficient estimates provided are given in units of ordered logits, or ordered log odds. Hence, for a one unit increase in the independent variable, the movement from the unvaccinated group to the vaccinated group is expected to change by its respective regression coefficient in the ordered log-odds scale while the other independent variables in the model are held constant. The ordinal logistic regressions were built using the statsmodel (version 0.14.2) library in Python and use the Broyden–Fletcher–Goldfarb–Shanno (BFGS) method for solving.
Qualitative
@@
To analyze our open-ended reasoning question, we take a multi-round, open-coding approach, similar to the approach used in horne2020tailoring . First, two of the authors individually went over the open-ended responses from the vaccinated group and the unvaccinated group to identify broad codes, continuing until the list of codes was stable. The two authors then discuss their codes to develop a combined list of codes, each with specific definitions of what reasoning fits into the category. The final set of codes can be found in Tables 6 and 7 at the end of the paper. Then the same two authors coded all of the responses for both groups (vaccinated and not). Since the respondents could describe as many reasons as they wanted, each participants response could fall into more than one category. After the second round of coding, agreement was computed, where a response is said to have agreement if all codes match across the two coders. In the unvaccinated group, 79.84% of the responses had full agreement, while the vaccinated group, 85.38% of the responses had full agreement. The two authors met again to discuss and resolve these conflicts. Disagreements in the unvaccinated group were mostly due to interpreting what entity was being distrusted (science, government, media, pharmaceutical companies, general), hence we merged these to a broader category of Distrust. For example, some response were very specific: “Because there’s not enough research to determine if it’s safe and I do not trust the government or media.” while others were more general: “i dont trust any of it”. Disagreements in the vaccinated group mostly were due to differentiating the categories Protect others versus Right thing to do. In this case, we opted for stricter definitions of each to avoid misinterpretation. For example, a response had to use the phrase “it was the right thing to do” to be included in the category Right thing to do, while calls to protecting others, such as “To protect my family, community” or “I wanted to protect others from catching Covid and help reduce the spread” were put in the category Protect others. Definitions and examples of the final agreed upon results can be found in Tables 5, 6, and 7.
@@ 3 Results
Figure 1: (a) Counties where at least one survey participant resides. Participants in our survey lived in 882 unique U.S. counties, with 471 counties represented in the vaccinated group and 461 counties represented in the unvaccinated group. (b) Correlation of key demographic variables used in study. Distributions of (c) political leaning, (d) rural-urban identity, and (e) education of vaccinated (blue) and unvaccinated (yellow) participants.
3.1 Demographic Differences